OpenStudy (anonymous):

How do I simplify (4+3i)(5-i) ? Step by step please!

OpenStudy (anonymous):

(4 +3i)(5-i) = 20 - 4i + 15i -3i^2 = 20 +11i +3 You would use FOIL to multiply all the terms. Combine like terms and since i^2 = -1, you can simplify.

OpenStudy (anonymous):

Thank you that really helped! Now I need to do a division problem, can you help me again? The problem is 3+i/5-i.